OpenStudy (anonymous):

how do you make a sketch of a square that has an area of 49x^2+84x+36?

OpenStudy (jim766):

factor it, each side will have a length of 1 factor

OpenStudy (mertsj):

Factor the area to find the length of each side.

OpenStudy (mertsj):

Draw a picture of a square and put the side length you found on the sides.

OpenStudy (jim766):

in this case 49x^2 + 84x + 36 = ( )( ) both factors will be the same

OpenStudy (anonymous):

(7x+6)(7x+6)

OpenStudy (jim766):

right.... or (7x =6)^2 7x + 6 is the length of each side (side)^2 = A of a square

OpenStudy (anonymous):

Thanks a million. This was very helpful. I just have to figure out how to complete the sketch.

OpenStudy (jim766):

its a sketch...doesnt have to be exact. draw a square and label each sides length